OpenStudy (anonymous):

Can someone help me with a couple questions? Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

OpenStudy (ccswims):

B is the only correct point-slope form

OpenStudy (ccswims):

I can help with more

OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6). 7 -14 0 14

OpenStudy (ccswims):

Using our rate of change formula we get -14, so B

OpenStudy (anonymous):

tysm! :) Choose the point-slope form of the equation below that represents the line that passes through the points (-6, 4) and (2, 0). y - 4 = -one half(x + 6) y - 4 = 2(x + 6) y + 6 = -one half(x - 4) y + 6 = 2(x - 4)

OpenStudy (anonymous):

I think it'd be A.

OpenStudy (ccswims):

Yea, its A

OpenStudy (ccswims):

Wait no, its C

OpenStudy (ccswims):

Wait no, I'm sorry it is A

OpenStudy (ccswims):

I just had like a brain fart, yea it's a

OpenStudy (ccswims):

sorry if that was confusing

OpenStudy (anonymous):

its okay!